Given the data below for the decomposition of NOBr (g) to NO (g) and Br2 (g), determine the order of the reaction with respect to NOBr (g).

To do this, create a graph of the data for each of the orders of reaction and report the R2 value for each linear fit. The graph with the R2 value closest to 1 is the most linear graph and the equation graphed corresponds to the order of the reaction.

*I do not need actual graphs made, just instructions on what to put into excel for x and y.

Time (s) [NOBr] (M) ln [NOBr] 1 / [NOBr] (M-1)
0 0.01 -4.60 100.000
2 0.0071 -4.95 140.845
4 0.0055 -5.20 181.818
6 0.0045 -5.40 222.222
8 0.0038 -5.57 263.158
10 0.0033 -5.71 303.030
Answer #1

Answer:

Given reaction is

2NOBr(g) ------> 2NO(g) + Br2(g)

Rate =k[NOBr]x

where k=rate constant, x=order with respect to NOBr.

From the given data,

The zero order plot is [NOBr] vs time.

The rate law is [NOBr]=-kt+[NOBr]0

                           y = mx + b

(Time on x-axis and [NOBr] on y-axis)

If the plot between [NOBr] vs time is straight line with good fitting, then the order of reaction is zero order and the slope of the line is equal to rate constant.

The first order plot is ln[NOBr] vs time.

The rate law is ln[NOBr]=-kt+ln[NOBr]0

                           y = mx + b

(Time on x-axis and ln[NOBr] on y-axis)

If the plot between ln[NOBr] vs time is straight line with good fitting, then the order of reaction is first order and the slope of the line is equal to rate constant.

The second order plot is [NOBr] vs time.

The rate law is 1/[NOBr]=kt+1/[NOBr]0

                           y = mx + b

(Time on x-axis and 1/[NOBr] on y-axis)

If the plot between 1/[NOBr] vs time is straight line with good fitting, then the order of reaction is second order and the slope of the line is equal to rate constant.

The plots are attached below

The zero order plot:

The first order plot:

The second order plot:

From the above plots, 1/[NOBr] vs time gives a straight line with R2=0.9999. Therefore the order of the reaction is second order. The rate constant, k=20.32 M-1 s-1.

Rate =(20.32 M-1 s-1)[NOBr]2

Please lat me know if you have any doubt. Thanks and i hope you like it.